OpenStudy (anonymous):

HELLO EVERYONE! Solve for x: -3(x + 3) = -3(x + 1) - 5. A)6 B)-6 C)All real numbers OR D)No solution

OpenStudy (emmamink):

let's distribute each side first.

OpenStudy (emmamink):

-3x + -9 = -3x + -3 - 5.

OpenStudy (anonymous):

-3(x+3) = -3(x+1)-5 -3x-9 = -3x-3-5 The -3's are distributed -3x-9 = -3x-8 It was simplified -9 = -8 The x's cancel out. The answer is no solution because the x's cancel out and -9 does not equal -8. So the answer is no solution.

OpenStudy (emmamink):

Now, do you know what the addition and multiplication property of equality are? It states that whatever you do to one side you must do to the other.

OpenStudy (emmamink):

Come on! Everyone always steals the answer right from under my feet and they get the medal instead.

OpenStudy (anonymous):

Thank you both! I'm sorry @EmmaMink for not giving you the medal! agupta's response came a couple seconds before yours & I gave it to him before I saw your post

OpenStudy (emmamink):

Oh.